# A Module-theoretic Introduction to Abstract Elementary Classes

### v1

@@
Abstract Elementary Classes have an entirely semantic definition, but one that is wide enough to encompass classes axiomatized by a whole host of different logics (see Example 2.5). Abstract Elementary Classes were first introduced in print in [She87], and there has been much work developing there classification theory. For references, see [Bal09, Gro1X, She09b, BV17b]. In the decades since this project began, the focus has mostly been on the abstract development of the theory: fixing an Abstract Elementary Class $\mathbb{K}$, assuming some additional structural properties (like amalgamation, see Section 2.5) and potentially some set-theoretic axioms, and developing classification theory.
@@
Section 2 is dedicated to this first goal. Understanding the strange of initially off-putting definition of Abstract Elementary Classes, Section 2.1 work to unpack the idea behind what makes a class an Abstract Elementary Classes by working through two basic (and hopefully instructive) examples: modules with submodule, and modules with pure submodule. Then, a great many examples are given, both from a model-theoretic/logical perspective (Example 2.5) and a module-theoretic perspective (Example 2.7). Finally, we introduce some of the basic notions and properties of Abstract Elementary Classes.
@@ 2.1. Motivation for definition
In each case, we describe a way of analyzing and understanding these groups that will be (mostly) typical of Abstract Elementary Classes. In the precise langauge we will define, both $(\mathbb{K}^{\text{R-Mod}},\subset)$ and $(\mathbb{K}^{\text{R-Mod}},\subset^{\text{pr}})$ are Abstract Elementary Classes with Löwenheim-Skolem number $\aleph_{0}+|R|$, while $(\mathbb{K}^{\text{R-Mod}},\subset)$ has the special property of being a universal class (see Definition 2.16.(1)). For those familar with category theory, the construction we describe is a hands on-version of taking a an accessible category $\mathcal{K}$ with directed colimits, and consider the representation given by the Yoneda embedding $\text{Hom}_{\mathcal{K}}(\mathcal{K}_{0},-)$, where $\mathcal{K}_{0}$ are the generators given by accessibility; this is a concrete way to prove that accessible categories are equivalent to those axiomatizable in $\mathbb{L}_{\infty,\infty}$. Crucially, the classes considered are finitely accessible, so this axiomatization is in $\mathbb{L}_{\infty,\omega}$ and the classes are closed under directed colimits.
@@ 2.1.1. Modules with embeddings
  • We are concerned with infinite structures. $R$ might be a finite ring, but each module $M$ we discuss is probably implicitly assumed to be infinite.
  • Coming from model theory, the maps between structures are probably (and implicitly) assumed to be monomorphisms/injections. This is becasue we want them to preserve the atomic formula ‘$x\neq y$.’
@@
we can generate a submodule $\langle\mathbb{a}\rangle_{M}\subset M$. In this way we can decompose $M$ into a directed union of finitely generated submodule:
@@
With these three properties, the correspondence goes the other way! Fix a directed poset $(I,<)$ and $M_{i}\in\mathbb{K}^{\text{R-Mod}}$ for $i\in I$ with the additional property [^3] This property will be implicitly assumed going forward. that $i<j$ from $I$ imply that $M_{i}\subset M_{j}$. This allows us to form the directed union of the system
@@
Given $f:M\to N$, we can build a system of embeddings between their decompositions: for each $\mathbb{a}=a_{1},\dots,a_{n}\in M$, define
$\bar{f}_{\mathbb{a}}:\langle a_{1},\dots,a_{n}\rangle_{M}$$\to$ $\langle f(a_{1}),\dots,f(a_{n})\rangle_{M}$
@@ 2.1.2. Pure embeddings
The contexts above does not quite provide the generality appropriate for Abstract Elementary Classes. Given a module $M$ and finite tuple $\mathbb{a}\in M$, there is a unique smallest submodule of $M$ containing $\mathbb{a}$, namely $\langle\mathbb{a}\rangle_{M}$. However, not every notion of subobjects that we want to study satisfies this property. Namely, this is not satisfied by the notion of pure submodule. There are many equivalent definitions of pure submodule, but the following is the most useful for our purpose.
@@
If $\exists\mathbb{x}\in{}^{n}N$ so $A\mathbb{x}=\mathbb{b}$,
then $\exists\mathbb{y}\in{}^{n}M$ so $A\mathbb{y}=\mathbb{b}$
We would like to tell the same story for $(\mathbb{K}^{\text{R-Mod}},\subset_{\text{pr}})$ that we did above, but the problem is that there are $\mathbb{a}\in M$ such that $\langle\mathbb{a}\rangle_{M}\not\subset_{\text{pr}}M$. For instance, taking $M=\mathbb{Z}/8\mathbb{Z}$ and $a=4$, $M$ thinks that $4$ is divisible by $2$ (expressed as a solution to `$[2][x_{1}]=[4]’$ ’). But $\langle 4\rangle_{M}=\{0,4\}$ does not. So we cannot find a generated pure submodule covering some parameters. However, the following means there is some pure cover of parameter sets.
@@
Note that more is possible: pure submodules are closed under intersection (see Definition 2.16.(1)), so there is a smallest pure submodule covering $X$. However, we suppress this additional structure to better represent the properties of general Abstract Elementary Classes.
This means that, given a $R$-module $M$, we can decompose it as a directed colimit of small pure submodules, just as before! There are two crucial changes here:
@@
    Beyond size, the new structures are no longer generated by anything. In the submodule case, a map between (free) generators $\mathbb{a}\in M\mapsto\mathbb{b}\in N$ lifts uniquely to a map between generated submodules $\langle\mathbb{a}\rangle_{M}\to\langle\mathbb{b}\rangle_{N}$. However, no such lift is possible with the pure submodules.
@@
This final point exactly follows from the failure of the small models to be generated by finite tuples. However, we can salvage this result by increasing the directedness of the systems of small models. Recall that for a cardinal $\kappa$, a partial order $(I,<)$ is $\kappa$-directed iff for all $X\subset I$ of size $<\kappa$, there is $i_{*}\in I$ so $i<i_{*}$ for all $i\in X$. If we upgrade the systems in Figure 2 to being $(|R|+\aleph_{0})^{+}$-directed, we can recover it. This construction parallels the $\mathbb{K}^{up}$ construction in [She09a, Lemma 1.24].
@@
    Moreover, one such system is given by setting $I=\mathcal{P}_{\lambda}M$ (the $<\lambda$-sized subsets of $M$) and, for $X\in I$, picking $M_{X}\subset_{\text{pr}}M$ to be a pure submodule containing
@@ 2.2. Basic definitions
    (Tarski-Vaught axiom) the class is closed under directed colimits which are created in Set;444‘Created in Set’ means that the underlying set of the colimits is the colimits of the underlying sets. A crucial contrast is with the category of Hilbert spaces: this category is closed under directed colimits, but the colimit adds more elements than exist in the constituent parts. Lieberman-Rosicky-Vasey [LRV23, Theorem 18] go farther and show that there is no way to make Hilbert spaces concrete so directed colimits are created in Set. Metric Abstract Elementary Classes (introduced by Hirvonen and Hyttinen [HH09]) are a formalism appropriate for (Cauchy) complete structures. and
@@
A strong embedding or $\mathbb{K}$-embedding $f:M\to N$ is a $\tau$-embedding so $f(M)\prec_{\mathbb{K}}N$. All maps that we write are assumed to be $\mathbb{K}$-embeddings.
@@ 2.3. Examples
A logic $\mathcal{L}$ then takes a language as input and outputs a set of formulas and axioms–more formally called sentences–that a structure can satisfy or not. When considering AECs, we also want to associate $\mathcal{L}$ with a strong substructure relation $\prec_{\mathcal{L}}^{*}$ that will turn an class axiomatized by a theory in $\mathcal{L}$ into an Abstract Elementary Class.
@@
    Fact 2.6 (Compactness Theorem, Gödel, Mal’cev).
@@
    For instance, in abelian groups, a formula $\phi(x)$ could only place finitely many restrictions on the order of $x$. But the logic $\mathbb{L}_{\omega_{1},\omega}$ can include the infinite information that it has no order in a single formula:
@@
      To fix this, for any negative instance $\neg Q_{\alpha}x\phi(x,\mathbb{y})$ in the fragment, we add the condition that, for $\mathbb{a}\in M$,
@@
      • The equivalence class quantifier [^9] I am not sure of the history of this quantifier, but I first saw it in [KMV21, Section 3]. $Q^{ec}_{\alpha}$, where
@@
        expresses that there is a set $X\subset M$ of size $\geq\aleph_{\alpha}$ such that any $n$-tuple $b_{1},\dots,b_{n}$ from $X$ satisfies $\phi(x_{1},\dots,x_{n};\mathbb{a})$. Imagining that $\phi$ assigns a coloring the the tuple explains the name Ramsey quantifier.
@@
      Although these notions might seem immediately connected to module theory, the, e.g., Magidor-Malitz quantifiers have been used in the model theory of modules [Bau84].
@@
      iff $\psi(x,y;\mathbb{a})$ defines a linear order on the set $\phi(M;\mathbb{a})$ with cofinality $\kappa$. Note that defining a linear is first-order expressible, so the strength really comes from the statement about the cofinality. The strong substructure to make Abstract Elementary Classes was given in [Bon23, Section 3]; it is technical so we don’t reproduce it here.
@@
Below, we list several natural algebraic classes that form Abstract Elementary Classes. Unless otherwise state, these classes form AECs using either the submodule relation or the pure submodule relation. To streamline presentation, we will list properties of these classes as we introduce them here, even though these properties are not introduced until later (mostly see Section 2.5). Most of the examples listed are modules and their variants.
@@
    Modules with purity: Another common class $(\mathbb{K}^{\text{R-Mod}},\subset^{\text{pr}})$, which uses pure substructure in place of substructure. Note the underlying models are the same, but we have changed the substructure.
    Although it has intersections instead of being universal, it still satisfies all of the nice properties listed for modules.
@@
    Finitely Butler groups: Recall that a torsion-free group is Butler iff it is a finite rank, pure subgroup of a completely decomposable group. Then $G$ is finitely Butler iff every finitely generated subgroup is a Butler group. The class $(\mathbb{K}^{fB},\subset_{\text{pr}})$ of. finitely Butler groups with pure embedding was first considered as an Abstract Elementary Class in [MA20, Section 5]. The proof that this is an AEC is formulaic as a locally verifiable property without unpacking what a Butler group is: if $\mathcal{B}$ is the class of all finitely generated Butler group with generators, then $\mathbb{K}^{fB}$ is axiomatized by the $\mathbb{L}_{|\mathcal{B}|^{+},\omega}$-sentence
@@
    ${}^{\perp}\mathcal{N}$ $=$ $\{M\in\mathbb{K}^{\text{R-Mod}}:\text{Ext}^{n}(M,N)=0\text{ for all }N\in\mathcal{N}\text{ and }n<\omega\}$
@@
    The classes $\left({}^{\perp}\mathcal{N},\prec_{\perp}\right)$ were first examined as an Abstract Elementary Class by Baldwin, Ekloff, and Trlifaj [BET07] and later in [Trl09, Trl08] and more. These classes represent an important gap in understanding how to characterize Abstract ELementary Classes: they are one of the very few classes where the strong substrcture relationship is not given syntactically and one of the very few classes where an axiomatization in one of the above logics is not known.
    This class is not always an Abstract Elementary Class, but Trlifaj [Trl08, Theorem 4.4] characterizes the Roots of Ext that are Abstract Elementary Classes as those ${}^{\perp}\mathcal{N}$ so ${}^{\perp}\mathcal{N}$ is deconstructible Definition 3.7.(2) and closed under direct limits; deconstructibility gives the Löwenheim-Skolem Axiom 2.4.(5) and the direct limits give the colimits propert of the Tarski-Vaught Axiom 2.4.(4).
@@
    Flat modules: The class $(\mathbb{K}^{flat},\subset_{\text{pr}})$ of flat modules form an Abstract ELementary Class. It has amalgamation, joint embedinng, no maximal models, and is tame [LRV, Section 6].

### v2

@@
Abstract.
The first-order model theory of modules has been studied for decades. More recently, the model theoretic study of nonelementary classes of modules–especially Abstract Elementary Classes of modules–has produced interesting results. This survey aims to discuss these recent results and give an introduction to the framework of Abstract Elementary Classes for module theorists.
1. Introduction
Model theory is the study of classes of structures that satisfy some fixed set of axioms. This makes it well-suited to interact with modules, as this is a class of structures that satisfy the module axioms. Indeed, there has been a rich history of interaction between model theory and module theory, for which [Pre88] is an excellent reference.
Most of this interaction has taken place in the context of first-order model theory. This is the most well-developed type of model theory, largely due to the Compactness Theorem (Fact 2.7) and other powerful properties of first-order logic. This is great as several classes of modules are axiomatizable in first-order logic: all modules, torsion-free modules, etc. To make use of the most powerful tools of model theory, one fixes a complete first-order theory of modules (given by specifying invariants conditions [Pre88, Corollary 2.13]) and looks at all modules over the fixed ring satisfying that theory. This allows one to make use of the powerful results of classification theory, first laid out by Shelah [She90].
If this was the end of the story, this article wouldn’t exist. On the module-theoretic side, there are many natural classes of modules that are not axiomatizable in first-order logic: locally pure injective modules, $\aleph_{1}$-free abelian groups, flat modules, etc. (see Example 2.8 for a long list). On the model theoretic-side, there are many classes of structures that one wants to understand and develop classfication theory for that are not elementary [^1] A class is elementary if it is the models of a complete first-order theory. A class is nonelementary if it is not elementary, normally only applied to a class of structures in a common language closed under isomorphisms.. One could study these classes case-by-case or even by looking at the logics that axiomatize them, but the most popular framework is that of Abstract Elementary Classes (Definition 2.5).
Abstract Elementary Classes have an entirely semantic definition, but one that is wide enough to encompass classes axiomatized by a whole host of different logics (see Example 2.6). Abstract Elementary Classes were introduced by Shelah and first appeared in print in Makowsky [BF85, Chapter XX, Definition 1.2.6.(ii)] (where they are called ‘abstract classes with Löwenheim number’). Their has been much work developing there classification theory; for references, see [Bal09, Gro1X, She09b, BV17b]. In the decades since this project began, the focus has mostly been on the abstract development of the theory: fixing an Abstract Elementary Class $\mathbb{K}$, assuming some additional structural properties (like amalgamation, see Section 2.5) and potentially some set-theoretic axioms, and developing classification theory.
@@ 1. Introduction
Section 2 is dedicated to this first goal. Recognizing the strange and technical definition of Abstract Elementary Classes, Section 2.1 works to unpack the idea behind what makes a class an Abstract Elementary Classes by working through two basic (and hopefully instructive) examples: modules with submodule, and modules with pure submodule. Then, a great many examples are given, both from a model-theoretic/logical perspective (Example 2.6) and a module-theoretic perspective (Example 2.8). Finally, we introduce some of the basic notions and properties of Abstract Elementary Classes.
@@
We would like to thank John Baldwin, Marcos Mazari-Armida, and the anonymous referee for helpful comments on an initial draft.
@@ 2.1. Motivation for definition
In each case, we describe a way of analyzing and understanding these cases that will be (mostly) typical of Abstract Elementary Classes. In the precise language we will provide, both $(\mathbb{K}^{\text{R-Mod}},\subset)$ and $(\mathbb{K}^{\text{R-Mod}},\subset^{\text{pr}})$ are Abstract Elementary Classes with Löwenheim-Skolem number $\aleph_{0}+|R|$, while $(\mathbb{K}^{\text{R-Mod}},\subset)$ has the special property of being a universal class (see Definition 2.18.(1)). For those familiar with category theory, the construction we describe is a hands on-version of taking an accessible category $\mathcal{K}$ with directed colimits, and consider the representation given by the Yoneda embedding $\text{Hom}_{\mathcal{K}}(\mathcal{K}_{0},-)$, where $\mathcal{K}_{0}$ are the generators given by accessibility; this is a concrete way to prove that accessible categories are equivalent to those classes that are axiomatizable in $\mathbb{L}_{\infty,\infty}$. Crucially, the classes considered are finitely accessible, so this axiomatization is in $\mathbb{L}_{\infty,\omega}$ and the classes are closed under directed colimits.
@@ 2.1.1. Modules with embeddings
  • We are concerned with infinite structures. $R$ might be a finite ring, but each module $M$ we discuss is implicitly assumed to be infinite.
  • Coming from model theory, the maps between structures are implicitly assumed to be monomorphisms/injections. This is becasue we want them to preserve the atomic formula ‘$x\neq y$.’
@@
we can generate a submodule $\langle\mathbb{a}\rangle_{M}\subset M$. In this way we can decompose $M$ into a directed union of all of its finitely generated submodules:
@@
With these three properties, the correspondence goes the other way! Fix a directed poset $(I,<)$ and finitely generated $M_{i}\in\mathbb{K}^{\text{R-Mod}}$ for $i\in I$ with the additional property [^3] This property will be implicitly assumed going forward. that $i<j$ from $I$ imply that $M_{i}\subset M_{j}$. This allows us to form the directed union of the system
@@
Given $f:M\to N$, we can build a system of embeddings between their decompositions: for each $\mathbb{a}=a_{1},\dots,a_{k}\in M$, define
$\bar{f}_{\mathbb{a}}:\langle a_{1},\dots,a_{n}\rangle_{M}$$\to$ $\langle f(a_{1}),\dots,f(a_{k})\rangle_{N}$
@@ 2.1.2. Pure embeddings
The context above does not quite provide the generality appropriate for Abstract Elementary Classes. Given a module $M$ and finite tuple $\mathbb{a}\in M$, there is a unique smallest submodule of $M$ containing $\mathbb{a}$, namely $\langle\mathbb{a}\rangle_{M}$. However, not every notion of strong subobject that we want to study satisfies this uniqueness property. Namely, this is not satisfied by the notion of pure submodule. There are many equivalent definitions of pure submodule, but the following is the most useful for our purpose.
@@
if $\exists\mathbb{x}\in{}^{n}N$ s.t. $A\mathbb{x}=\mathbb{b}$,
then $\exists\mathbb{y}\in{}^{n}M$ s.t. $A\mathbb{y}=\mathbb{b}$
We would like to tell the same story for $(\mathbb{K}^{\text{R-Mod}},\subset_{\text{pr}})$ that we did above, but the problem is that there are $\mathbb{a}\in M$ such that $\langle\mathbb{a}\rangle_{M}\not\subset_{\text{pr}}M$. For instance, taking $M=\mathbb{Z}/8\mathbb{Z}$ and $a=4$, $M$ thinks that $4$ is divisible by $2$ (expressed as a solution to `$[2][x_{1}]=[4]’$ ’). But $\langle 4\rangle_{M}=\{0,4\}$ does not. So we cannot find a generated pure submodule covering some parameters. However, the following means there is some pure cover of any given parameter set.
@@
In Definition 2.18.(1), a weakening of universality called closure under intersections is introduced (but it’s mostly what it sounds like). The class $(\mathbb{K}^{\text{R-Mod}},\subset_{\text{pr}})$ is not closed under intersections; thanks to the referee for providing this example.
Example 2.3.
We work with abelian groups (so $R=\mathbb{Z}$). Set $N=\mathbb{Z}(2^{\infty})$ to be the [^4] This example works with any prime $p$. Prüfer 2-group with generators $g_{n}$ so $2g_{n+1}=g_{n}$ and $02g_{0}=$ 0.
Define the modules
$M$ $=$ $N\otimes N$
$M_{0}$ $=$ $N\otimes\{0\}$
$M_{1}$ $=$ $\langle(g_{n},2g_{n}):n<\omega\rangle$
Then $M_{0},M_{1}$ are both pure submodules of $M$, as witnessed by the fact that direct summing with $\{0\}\otimes N$ gives $M$. Then the intersection
$M_{0}\cap M_{1}=\left\{(g_{0},0),(0,0)\right\}$
is not a pure submodule of $M$ since $(g_{0},0)$ is not divisible by 2.
Lemma 2.2 means that, given a $R$-module $M$, we can decompose it as a directed colimit of small pure submodules, just as before! There are two crucial changes here:
@@
    The new structures are no longer generated by anything. In the submodule case, a map between (free) generators $\mathbb{a}\in M\mapsto\mathbb{b}\in N$ lifts uniquely to a map between generated submodules $\langle\mathbb{a}\rangle_{M}\to\langle\mathbb{b}\rangle_{N}$. However, no such lift is possible with the pure submodules.
@@
[…]
@@
    Moreover, one such system is given by setting $I=\mathcal{P}_{\lambda}(M)$ (the $<\lambda$-sized subsets of $M$) and, for $X\in I$, picking $M_{X}\subset_{\text{pr}}M$ to be a pure submodule containing
@@ 2.2. Basic definitions
    (Tarski-Vaught axiom) the class is closed under directed colimits which are created in Set; [^5] The traditional phrasing of this axiom is about closure under unions of chains. However, a classic result (see [MKS64, Theorem 2]) says that these are equivalent conditions.
    ‘Created in Set’ means that the underlying set of the colimits is the colimit of the underlying sets. A crucial contrast is with the category of Hilbert spaces: this category is closed under directed colimits, but the colimit adds more elements than exist in the constituent parts. Lieberman-Rosicky-Vasey [LRV23, Theorem 18] go farther and show that there is no way to make Hilbert spaces concrete so directed colimits are created in Set. Metric Abstract Elementary Classes (introduced by Hirvonen and Hyttinen [HH09]) are a formalism appropriate for (Cauchy) complete structures.
    and
@@
$\mathbb{K}_{<\lambda}$ $=$ $\{M\in\mathbb{K}:\|M\|<\lambda\}$
$\mathbb{K}_{\geq\lambda}$ $=$ $\{M\in\mathbb{K}:\|M\|\geq\lambda\}$
@@
Note that, due to the Tarski-Vaught axiom, $\mathbb{K}_{\lambda}$ uniquely determines $\mathbb{K}_{\geq\lambda}$ by taking the closure under directed colimits.
[…]
@@ 2.3. Examples
A logic $\mathcal{L}$ then takes a language as input and outputs a set of formulas and axioms–more formally called sentences–that a structure can satisfy or not. When considering AECs, we also want to associate $\mathcal{L}$ with a strong substructure relation $\prec_{\mathcal{L}}^{*}$ that will turn a class axiomatized by a theory in $\mathcal{L}$ into an Abstract Elementary Class.
@@
    Fact 2.7 (Compactness Theorem, Gödel, Mal’cev).
@@
    For instance, in abelian groups, a formula $\phi(x)$ could only place finitely many restrictions on the order of $x$. But the logic $\mathbb{L}_{\omega_{1},\omega}$ can include the infinite information that an element is not of finite order in a single formula:
@@
      To fix this, when we write ‘$M\prec_{\mathbb{L}(Q)}^{*}N$, we add the condition that, for any negative instance $\neg Q_{\alpha}x\phi(x,\mathbb{y})$ in the fragment and $\mathbb{a}\in M$, we have
@@
      • The equivalence class quantifier [^10] This and many other quantifiers are discussed in [BF85, Chapters IV and VI]. $Q^{ec}_{\alpha}$, where
@@
        expresses that there is a set $X\subset M$ of size $\geq\aleph_{\alpha}$ such that any $n$-tuple $b_{1},\dots,b_{n}$ from $X$ satisfies $\phi(x_{1},\dots,x_{n};\mathbb{a})$. Imagining that $\phi$ assigns a coloring the tuple explains the name Ramsey quantifier.
@@
      Although these notions might not seem immediately connected to module theory, the, e.g., Magidor-Malitz quantifiers have been used in the model theory of modules [Bau84].
@@
      iff $\psi(x,y;\mathbb{a})$ defines a linear order on the set $\phi(M;\mathbb{a})$ with cofinality $\kappa$. Note that defining a linear order is first-order expressible, so the strength really comes from the statement about the cofinality. The strong substructure to make Abstract Elementary Classes was given in [Bon23, Section 3]; it is technical so we don’t reproduce it here.
@@
Below, we list several natural algebraic classes that form Abstract Elementary Classes. Unless otherwise stated, these classes form AECs using either the submodule relation or the pure submodule relation. To streamline presentation, we will list properties of these classes as we introduce them here, even though these properties are not introduced until later (mostly see Section 2.5). Most of the examples listed are modules and their variants.
@@
    Modules with purity: Another common class $(\mathbb{K}^{\text{R-Mod}},\subset^{\text{pr}})$, which uses pure substructure in place of substructure. Note the underlying models are the same, but we have changed the notion of strong substructure.
    It is not universal (or closed under intersections), but it still satisfies all of the nice properties listed for modules.
@@
    Finitely Butler groups: Recall that a torsion-free group is Butler iff it is a finite rank, pure subgroup of a completely decomposable group. Then $G$ is finitely Butler iff every finitely generated subgroup is a Butler group. The class $(\mathbb{K}^{fB},\subset_{\text{pr}})$ of finitely Butler groups with pure embedding was first considered as an Abstract Elementary Class in [MA20, Section 5]. The proof that this is an AEC is formulaic as a locally verifiable property without unpacking what a Butler group is: if $\mathcal{B}$ is the class of all finitely generated Butler group with generators, then $\mathbb{K}^{fB}$ is axiomatized by the $\mathbb{L}_{|\mathcal{B}|^{+},\omega}$-sentence
@@
    ${}^{\perp}\mathcal{N}$ $=$ $\{M\in\mathbb{K}^{\text{R-Mod}}:\text{Ext}^{n}(M,N)=0\text{ for all }N\in\mathcal{N}\text{ and }0<n<\omega\}$
@@
    The classes $\left({}^{\perp}\mathcal{N},\prec_{\perp}\right)$ were first examined as an Abstract Elementary Class by Baldwin, Eklof, and Trlifaj [BET07] and later in [Trl09, Trl08] and more. These classes represent an important gap in understanding how to characterize Abstract Elementary Classes: they are one of the very few classes where the strong substrcture relationship is not given syntactically and one of the very few classes where an axiomatization in one of the above logics is not known.
    The class $\left({}^{\perp}\mathcal{N},\prec_{\perp}\right)$ is not always an Abstract Elementary Class, but Trlifaj [Trl08, Theorem 4.4] characterizes the Roots of Ext that are Abstract Elementary Classes as those ${}^{\perp}\mathcal{N}$ so ${}^{\perp}\mathcal{N}$ is deconstructible (Definition 3.7.(2) and closed under direct limits; deconstructibility gives the Löwenheim-Skolem Axiom 2.5.(5) and the direct limits give the colimits propert of the Tarski-Vaught Axiom 2.5.(4).
@@
    Flat modules: The class $(\mathbb{K}^{flat},\subset_{\text{pr}})$ of flat modules form an Abstract Elementary Class. It has amalgamation, joint embedinng, no maximal models, and is tame [LRV, Section 6]. If $R$ is a Prüfer domain, then this class is closed under intersections.